OpenStudy (anonymous):

Let f(x) = 4x - 5 and g(x) = 6x - 3. Find f(x) - g(x).

OpenStudy (johnweldon1993):

So just do 4x - 5 - 6x - 3 then combine like terms

OpenStudy (anonymous):



OpenStudy (anonymous):

okay i understand thank you

OpenStudy (anonymous):

4x -5 - (6x -3) =0 4x - 5 - 6x + 3 = 0 -2x - 2 =0 -2x=2 X= -1